OpenStudy (anonymous):

f(x)=(x^2-3x-4)/(x-2) Find points of increase and decrease and all relative extrema.

OpenStudy (anonymous):

you don't really need any calculus for this one

OpenStudy (anonymous):

Try factoring the numerator

OpenStudy (anonymous):

it is a rational function with a slant asymptote at \(x-1\)

OpenStudy (anonymous):

it is always increasing

OpenStudy (anonymous):

you could take the derivative and get \[\frac{x^2-4x+10}{(x-2)^2}\] but the denominator is never negative, and neither is the numerator ( you can check that it has no zeros)

OpenStudy (anonymous):

if you remember plotting rational functions in pre calc you may remember what something like this looks like

OpenStudy (anonymous):

That's the derivative that I got also. I was trying to find the critical points from this...

OpenStudy (anonymous):

So what would the critical points be?

OpenStudy (anonymous):

Factorise the Numerator... Can you??? Then the critical points for an expression like this... \[(x-a)(x-b)/ (x-c) \] are a,b and c...

OpenStudy (anonymous):

I can't factor \[x^2-4x-10\]

OpenStudy (anonymous):

You are supposed to factor x^2 - 3x-4 ... I think your question says so...

OpenStudy (anonymous):

Into... (x-2)(x-2)...?

OpenStudy (anonymous):

x^2 - 3x - 4 =x^2 - 4x +x -4 = x(x-4) +1(x-4) = (x-4) (x+1)

OpenStudy (anonymous):

And critical points are then x=4 and x=-1

OpenStudy (anonymous):

I have to find the first derivative then factor